# Long distance propagation of wave beams in paraxial regime

### v1

@@ 4.3 Proof of the tightness result
$=\mathbb{E}|\phi^{\varepsilon}(z,r,x+h)-\phi^{\varepsilon}(z,r,x)|^{2n}:=\mathbb{E}|\delta_{h}\phi^{\varepsilon}(z,r,x)|^{2n}\,,$
@@
where $0<\alpha\leq 1$ in the kinetic regime and $10<\alpha<$ 1 in the diffusive regime.
@@
$|h|^{2n}\theta\eta^{2n-2}e^{\frac{\lambda[R,n,n,z,1]}{\eta^{2}}}\|\prod\limits_{m=1}^{2n}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|,$
where from (36), $\lambda[R,n,n,z,1]=c(n,z)\|\langle k\rangle^{2+2n}\hat{R}(k)\|$. Due to assumption (9), we can find an $n$ such that $12n\geq d+$ 1 and $\|\langle k\rangle^{2+2n}\hat{R}(k)\|$ is bounded so that the term above is bounded by $|h|^{2n}$ (and thereby $|h|^{2n\alpha}$) uniformly in $\theta$.
@@ 5 Conclusions
Starting from the paraxial approximation, we have characterized the limiting statistical distribution of high frequency wavefields propagating over long distances through weak turbulence. In the scintillation regime considered here, we show that an appropriately rescaled macroscopic wavefield is asymptotically described by a complex Gaussian distribution with scintillation saturating to unity. This is consistent with experimental observations of speckle patterns where intensity of the wavefield is modeled by an exponential law. We have used a Gaussian assumption with short range correlations on the refractive index fluctuations to simplify our derivations. This allows us to show that the statistical moments of the wavefield of the paraxial wave equation are well approximated by moments given by closed-form PDEs of the Itô-Schrödinger model. Our analysis is primarily done in the Fourier domain with errors controlled in the total variation norm which translates to error estimates in the uniform sense in the physical domain. This choice is important for applications as physical manifestations like speckle are observed predominantly in the spatial domain. This in turn allows us to use the results in [5] where the moments of the Itô-Schrödinger equation have been shown to converge to that of a complex Gaussian random variable. The extension of these results to a more general class of turbulence scenarios where the Gaussian assumption on the randomness is relaxed is left for future investigation.
@@ Acknowledgments
We would like to thank Christophe Gomez for many stimulating discussions on the validity of the paraxial and Itô-Schrödinger models. This work was funded in part by NSF grant DMS-230641.

### v2

@@ 4.3 Proof of the tightness result
$=\mathbb{E}|\phi^{\varepsilon}(z,r,x+h)-\phi^{\varepsilon}(z,r,x)|^{2n}=\mathbb{E}|\delta_{h}\phi^{\varepsilon}(z,r,x)|^{2n}\,$,
@@
where $0<\alpha\leq 1$ in the kinetic regime and $0<\alpha<1$ in the diffusive regime.
@@
$|h|^{2n}\theta\eta^{2n-2}e^{\frac{\lambda[{\color[rgb]{0,0,0}C},n,n,z,1]}{\eta^{2}}}\|\prod\limits_{m=1}^{2n}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|,$
where from (36), $\color[rgb]{0,0,0}\lambda[C,n,n,z,1]=c(n,z)\int_{\mathbb{R}^{d+1}}\langle s\rangle\langle k\rangle^{2+2n}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k$}. Due to assumption (9), we can find an $n$ such that $2n\geq d+1$ and $\|\langle k\rangle^{2+2n}\hat{R}(k)\|$ is bounded so that the term above is bounded by $|h|^{2n}$ (and thereby $|h|^{2n\alpha}$) uniformly in $\theta$.
@@
4.4 Proof of Theorem 3.6
Proof.
It suffices to verify that
$m^{\theta}_{p_{1}+p_{2},q_{1}+q_{2}}(z,r,X,Y,r^{\prime},X^{\prime},Y^{\prime})=\mathbb{E}\prod_{j=1}^{p_{1}}\phi^{\theta}(z,r,x_{j})\prod_{j=1}^{p_{2}}\phi^{\theta}(z,r^{\prime},x^{\prime}_{j})\prod_{l=1}^{q_{1}}\phi^{\theta\ast}(z,r,y_{l})\prod_{l=1}^{q_{2}}\phi^{\theta\ast}(z,r^{\prime},y^{\prime}_{l})$
is asymptotically equal to
$m_{p_{1},q_{1}}(z,r,X,Y)m_{p_{2},q_{2}}(z,r^{\prime},X^{\prime},Y^{\prime})=\mathbb{E}[\prod_{j=1}^{p_{1}}\phi(z,r,x_{j})\prod_{l=1}^{q_{1}}\phi^{\ast}(z,r,y_{l})]\mathbb{E}[\prod_{j=1}^{p_{2}}\phi(z,r^{\prime},x^{\prime}_{j})\prod_{l=1}^{q_{2}}\phi^{\ast}(z,r^{\prime},y^{\prime}_{l})]\,.$
From Theorem 4.6, $m^{\theta}_{p_{1}+p_{2},q_{1}+q_{2}}$ is well approximated by sums of products of second moments. We have that this is equal to the product $m_{p_{1},q_{1}}(z,r,X,Y)m_{p_{2},q_{2}}(z,r^{\prime},X^{\prime},Y^{\prime})$ if the two-point correlation between points centred at $r$ and $r^{\prime}$ are asymptotically zero. For simplicity, we only prove the case for when $\beta=1$. From Theorem 4.6,
$\mathbb{E}\phi^{\theta}(z,r,x)\phi^{\theta\ast}(z,r^{\prime},y^{\prime})=m_{1,1}^{\theta}(z,r,r^{\prime},x,y^{\prime})+c(z)(\varepsilon^{\frac{1}{3}}+\theta^{\frac{1}{2}})\,$,
where the second moment $m_{1,1}^{\theta}$ is given explicitly by [5]
$m_{1,1}^{\theta}(z,r,r^{\prime},x,y^{\prime})$ $=\int\limits_{\mathbb{R}^{2d}}e^{i\varepsilon^{-1}(\xi\cdot r-\zeta\cdot r^{\prime})}e^{i\eta(\xi\cdot x-\zeta\cdot y^{\prime})}e^{-\frac{i\eta z}{\varepsilon}(|\xi|^{2}-|\zeta|^{2})}\hat{u}^{\varepsilon}_{0}(\xi)\hat{u}^{\varepsilon\ast}_{0}(\zeta)$
$\times\exp\Big{(}\frac{z}{\eta^{2}}\int_{0}^{1}Q\big{(}\frac{r^{\prime}-r}{\varepsilon}+\eta(y^{\prime}-x)+\frac{2\eta sz}{\varepsilon}(\xi-\zeta)\big{)}\mathrm{d}s\Big{)}\frac{\mathrm{d}\xi\mathrm{d}\zeta}{(2\pi)^{2d}}\,$,
where $Q(x)=R(x)-R(0)\leq 0$. From a change of variables $(\xi/\varepsilon,\zeta/\varepsilon)\to(\xi,\zeta)$,
$m_{1,1}^{\theta}(z,r,r^{\prime},x,y^{\prime})$ $=\int\limits_{\mathbb{R}^{2d}}e^{i(\xi\cdot r-\zeta\cdot r^{\prime})}e^{i\eta\varepsilon(\xi\cdot x-\zeta\cdot y^{\prime})}e^{-{i\eta\varepsilon z}(|\xi|^{2}-|\zeta|^{2})}\hat{u}_{0}(\xi)\hat{u}^{\ast}_{0}(\zeta)$
$\times\exp\Big{(}\frac{z}{\eta^{2}}\int_{0}^{1}Q\big{(}\frac{r^{\prime}-r}{\varepsilon}+\eta(y^{\prime}-x)+{2\eta sz}(\xi-\zeta)\big{)}\mathrm{d}s\Big{)}\frac{\mathrm{d}\xi\mathrm{d}\zeta}{(2\pi)^{2d}}\,.$
Noting that
$\lim_{\varepsilon\to 0}\exp\Big{(}\frac{z}{\eta^{2}}\int_{0}^{1}Q\big{(}\frac{r^{\prime}-r}{\varepsilon}+\eta(y^{\prime}-x)+{2\eta sz}(\xi-\zeta)\big{)}\mathrm{d}s\Big{)}\to 0$
pointwise, the proof follows from Lebesgue’s dominated convergence theorem. ∎
@@ 5 Conclusions
Starting from the paraxial approximation, we have characterized the limiting statistical distribution of high frequency wavefields propagating over long distances through weak turbulence. In the scintillation regime considered here, we show that an appropriately rescaled macroscopic wavefield is asymptotically described by a complex Gaussian distribution with scintillation saturating to unity. This is consistent with experimental observations of speckle patterns where intensity of the wavefield is modeled by an exponential law. We have used a Gaussian assumption with short range correlations on the refractive index fluctuations to simplify our derivations. This allows us to show that the statistical moments of the wavefield of the paraxial wave equation are well approximated by moments given by closed-form PDEs of the Itô-Schrödinger model. Our analysis is primarily done in the Fourier domain with errors controlled in the total variation norm which translates to error estimates in the uniform sense in the physical domain. This choice is important for applications as physical manifestations like speckle are observed predominantly in the spatial domain. This in turn allows us to use the results in [5] where the moments of the Itô-Schrödinger equation have been shown to converge to that of a complex Gaussian random variable. The extension of these results to a more general class of turbulence scenarios where the Gaussian assumption on the randomness is relaxed and numerical simulations in the scintillation regime, say using operator splitting techniques [7] are left for future investigation.
@@ Acknowledgments
We would like to thank Christophe Gomez for many stimulating discussions on the validity of the paraxial and Itô-Schrödinger models. We also thank the two anonymous referees for their feedback in helping relax the assumptions on the random potential, suggesting the inclusion of Theorem 3.6 and in improving the overall presentation of this paper.